OpenStudy (anonymous):

A bicycle wheel has diameter twenty-six inches. Approximately how far does the bike travel when the wheel makes 5 complete rotations? eighty-one point six four inches one hundred sixty-three point two eight inches four hundred eight point two inches eight hundred sixteen point four inches

OpenStudy (anonymous):

i am giving out a medal to whoever answers it first

OpenStudy (anonymous):

please help me

OpenStudy (hoblos):

by each rotation the bicycle move a distance equal to the perimeter of the wheel the perimeter of the circle is \[2\pi r\]

OpenStudy (anonymous):

so that would be 14 then what would i do

OpenStudy (anonymous):

no 14 or 13 which one

OpenStudy (hoblos):

r= d/2 = 26/2 = 13 2*pi*r = 81.64 the bicycle moves 81.64 inches for each rotation for 5 rotations how much would it be ?

OpenStudy (anonymous):

ok thanks that helped a lot

OpenStudy (hoblos):

you got the answer ?

OpenStudy (anonymous):

yes 408.2

OpenStudy (hoblos):

correct.. welcome :)

OpenStudy (anonymous):

one last thing on this photo what is it called

OpenStudy (anonymous):